# Efficient parallel scheduling for sparse triangular solvers

### v1

@@ 8. Conclusion and future directions
The results show that our barrier list schedulers indeed significantly speed up the parallel SpTRSV kernel, reducing the execution time by a 1.45×1.45$\times$1.45 × geometric-mean factor compared to SpMP and 3.24×3.24$\times$3.24 × compared to HDagg on the SuiteSparse benchmark. The data also shows that the three main components of our approach (algorithms, coarsening, and reordering) indeed all contribute to the speed-up, and that the improvement is consistent over various matrix types and architectures.
Future work may consider the adaptation our algorithms to non-uniform memory access (NUMA) architectures. In particular, the AMD x86 data in Section 7.6 confirms that our algorithms scale well to a high number of cores. However, when solving SpTRSV on highly NUMA architectures, we expect our parallel execution schedules to be less effective. It is an interesting question whether one can develop scheduling algorithms that can also efficiently adapt to NUMA, for example, by considering non-uniform bandwidth or latency.
Another promising direction for future work is to combine our barrier list scheduling algorithms with other approaches that proved successful for SpTRSV in the past. For instance, one could seek to merge our algorithms with the block decomposition techniques described in Section 1.2, or to adapt them to a semi-asynchronous setting as in SpMP, in order to allow for a more flexible parallel execution. These methods could allow for further speed-ups on top of our current results.
@@ Appendix A Tables of matrices
We also include some basic statistics of the matrices used in the experiments, cf. Section 6.2.
@@
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] used for the evaluation. The average wavefront size has been rounded down., label = table:florida-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0, Matrix Dimension #Non-zeros Avg. wavefront
@@
[ theme = ams-theme, caption = Matrices and statistics in the Erdős–Rényi data set used for the evaluation. The average wavefront size has been rounded down., label = table:erdos-renyi-graph-table, ] colspec = l | r | r | r , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeroes Avg. wavefront
ErdosRenyi_100k_19m_A 100,000 19,999,021 109
ErdosRenyi_100k_19m_B 100,000 19,998,182 109
ErdosRenyi_100k_19m_C 100,000 19,997,897 107
ErdosRenyi_100k_19m_D 100,000 19,995,405 106
ErdosRenyi_100k_19m_E 100,000 19,994,516 107
ErdosRenyi_100k_19m_G 100,000 19,989,535 106
ErdosRenyi_100k_19m_H 100,000 19,999,989 110
ErdosRenyi_100k_1m_A 100,000 1,001,528 1,785
ErdosRenyi_100k_1m_B 100,000 1,000,452 1,818
ErdosRenyi_100k_1m_C 100,000 1,000,315 1,818
ErdosRenyi_100k_1m_E 100,000 1,000,044 1,666
ErdosRenyi_100k_1m_F 100,000 1,000,406 1,785
ErdosRenyi_100k_1m_G 100,000 1,001,171 1,724
ErdosRenyi_100k_1m_H 100,000 1,001,551 1,886
ErdosRenyi_100k_1m_I 100,000 1,000,237 1,639
ErdosRenyi_100k_1m_J 100,000 1,001,533 1,851
ErdosRenyi_100k_20m_F 100,000 20,001,732 107
ErdosRenyi_100k_20m_I 100,000 20,006,442 109
ErdosRenyi_100k_20m_J 100,000 20,003,479 109
ErdosRenyi_100k_4m_A 100,000 4,998,205 395
ErdosRenyi_100k_4m_C 100,000 4,999,271 398
ErdosRenyi_100k_4m_G 100,000 4,999,358 401
ErdosRenyi_100k_4m_J 100,000 4,996,501 414
ErdosRenyi_100k_5m_B 100,000 5,006,107 411
ErdosRenyi_100k_5m_D 100,000 5,001,575 404
ErdosRenyi_100k_5m_E 100,000 5,004,251 400
ErdosRenyi_100k_5m_F 100,000 5,002,190 400
ErdosRenyi_100k_5m_H 100,000 5,000,573 409
ErdosRenyi_100k_5m_I 100,000 5,001,846 400
ErdosRenyi_100k_999k_D 100,000 999,915 1,818
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics in the narrow bandwidth data set used for the evaluation. The average wavefront size has been rounded down., label = table:random-bandwidth-graph-table, ] colspec = l | r | r | r , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeroes Avg. wavefront
RandomBand_p14_b10_100k_146k_A 100,000 146,565 87
RandomBand_p14_b10_100k_146k_B 100,000 146,328 115
RandomBand_p14_b10_100k_146k_D 100,000 146,972 73
RandomBand_p14_b10_100k_146k_F 100,000 146,855 111
RandomBand_p14_b10_100k_146k_J 100,000 146,781 105
RandomBand_p14_b10_100k_147k_C 100,000 147,201 61
RandomBand_p14_b10_100k_147k_E 100,000 147,369 73
RandomBand_p14_b10_100k_147k_G 100,000 147,350 132
RandomBand_p14_b10_100k_147k_H 100,000 147,412 85
RandomBand_p14_b10_100k_147k_I 100,000 147,132 132
RandomBand_p3_b42_100k_127k_A 100,000 127,045 46
RandomBand_p3_b42_100k_127k_B 100,000 127,019 55
RandomBand_p3_b42_100k_127k_C 100,000 127,708 29
RandomBand_p3_b42_100k_127k_D 100,000 127,341 45
RandomBand_p3_b42_100k_127k_E 100,000 127,569 67
RandomBand_p3_b42_100k_127k_F 100,000 127,137 47
RandomBand_p3_b42_100k_127k_G 100,000 127,774 52
RandomBand_p3_b42_100k_127k_H 100,000 127,029 46
RandomBand_p3_b42_100k_127k_I 100,000 127,475 39
RandomBand_p3_b42_100k_127k_J 100,000 127,275 62
RandomBand_p5_b20_100k_102k_A 100,000 102,053 1,298
RandomBand_p5_b20_100k_102k_B 100,000 102,621 1,063
RandomBand_p5_b20_100k_102k_C 100,000 102,021 1,298
RandomBand_p5_b20_100k_102k_D 100,000 102,968 1,075
RandomBand_p5_b20_100k_102k_E 100,000 102,650 952
RandomBand_p5_b20_100k_102k_F 100,000 102,309 1,162
RandomBand_p5_b20_100k_102k_H 100,000 102,324 1,190
RandomBand_p5_b20_100k_102k_I 100,000 102,465 1,369
RandomBand_p5_b20_100k_102k_J 100,000 102,244 1,010
RandomBand_p5_b20_100k_103k_G 100,000 103,152 892
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] symmetrically permuted using the fill-reducing method ‘METIS_NodeND’ of [KK98]. The average wavefront size has been rounded down., label = table:hdagg-metis-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Dimension #Non-zeros Avg. wavefront

### v2

@@ 8. Conclusion and future directions
The results show that our GrowLocal scheduler indeed significantly speeds up the parallel SpTRSV kernel, reducing the execution time by a 1.42×1.42$\times$1.42 × geometric-mean factor compared to SpMP and 3.32×3.32$\times$3.32 × compared to HDagg on the SuiteSparse benchmark. The data also shows that the algorithm performs similarly well on multiple other data sets, and the improvements are consistent over various architectures. The scheduling time of GrowLocal is also competitive with the baselines, especially when combined with the block decomposition technique.
Future work may consider the adaptation of our algorithm to non-uniform memory access (NUMA) architectures. In particular, the AMD x86 data in Section 7.6 confirms that our algorithm scales well to a high number of cores. However, when solving SpTRSV on highly NUMA architectures, we expect the parallel execution to be less effective. In order to adapt to such a NUMA setting, one should consider fundamental changes to the SpTRSV kernel and the data structures, such as partitioning the matrix local to threads, interleaving the vector, and reducing the need for global synchronization. It is also an interesting question whether the scheduling algorithm can be efficiently adapted to NUMA, for example, by considering non-uniform bandwidth or latency. To our knowledge, there are currently no scheduling algorithms for SpTRSV that directly account for such NUMA effects.
Another promising direction for future work is to combine our barrier list scheduling algorithms with other approaches that proved successful for SpTRSV in the past. For instance, one could seek to adapt GrowLocal to a semi-asynchronous setting as in SpMP, in order to allow for a more flexible parallel execution. This could allow for further speed-ups on top of our current results.
@@ Appendix A Tables of matrices
Here we provide some basic statistics of the matrices used in the experiments, cf. Section 6.2.
@@
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] used for the evaluation. The average wavefront size (Avg. wf) has been rounded down., label = table:florida-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0, Matrix Size #Non-zeros Avg. wf
@@
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] symmetrically permuted using the fill-reducing method ‘METIS_NodeND’ of [KK98]. The average wavefront size (Avg. wf) has been rounded down., label = table:hdagg-metis-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeros Avg. wf
@@
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics from SuiteSparse Matrix Collection [DH11] post Eigen incomplete Cholesky [GJ+10] used for the evaluation. The average wavefront size (Avg. wf) has been rounded down., label = table:cholesky-graph-table, ] colspec = Q[l,m] | Q[r,m] | Q[r,m] | Q[r,m] , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeros Avg. wf
af_0_k101_iCh 503,625 9,027,150 195
af_shell7_iCh 504,855 9,046,865 668
apache2_iCh 715,176 2,766,523 79,464
audikw_1_iCh 943,695 39,297,771 138
bmw7st_1_iCh 141,347 3,740,507 340
bmwcra_1_iCh 148,770 5,396,386 89
bone010_iCh 986,703 36,326,514 340
boneS01_iCh 127,224 3,421,188 245
boneS10_iCh 914,898 28,191,660 521
Bump_2911_iCh 2,911,419 65,320,659 1,048
consph_iCh 83,334 3,046,907 78
Dubcova3_iCh 146,689 1,891,669 1,594
ecology2_iCh 999,999 2,997,995 142,857
Emilia_923_iCh 923,136 20,964,171 511
Fault_639_iCh 638,802 14,626,683 422
Flan_1565_iCh 1,564,794 59,485,419 689
G3_circuit_iCh 1,585,478 4,623,152 88,082
Geo_1438_iCh 1,437,960 32,297,325 768
hood_iCh 220,542 5,494,489 1,050
Hook_1498_iCh 1,498,023 31,207,734 649
inline_1_iCh 503,712 18,660,027 679
ldoor_iCh 952,203 23,737,339 3,317
msdoor_iCh 415,863 10,328,399 956
offshore_iCh 259,789 2,251,231 1,114
parabolic_fem_iCh 525,825 2,100,225 19,475
PFlow_742_iCh 742,793 18,940,627 240
Queen_4147_iCh 4,147,110 166,823,197 719
s3dkt3m2_iCh 90,449 1,921,955 104
Serena_iCh 1,391,349 32,961,525 940
shipsec1_iCh 140,874 3,977,139 259
StocF-1465_iCh 1,465,137 11,235,263 2,990
thermal2_iCh 1,228,045 4,904,179 47,232
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics in the Erdős–Rényi data set used for the evaluation. The average wavefront size (Avg. wf) has been rounded down., label = table:erdos-renyi-graph-table, ] colspec = l | r | r | r , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeroes Avg. wf
ER_100k_19m_A 100,000 19,999,021 109
ER_100k_19m_B 100,000 19,998,182 109
ER_100k_19m_C 100,000 19,997,897 107
ER_100k_19m_D 100,000 19,995,405 106
ER_100k_19m_E 100,000 19,994,516 107
ER_100k_19m_G 100,000 19,989,535 106
ER_100k_19m_H 100,000 19,999,989 110
ER_100k_1m_A 100,000 1,001,528 1,785
ER_100k_1m_B 100,000 1,000,452 1,818
ER_100k_1m_C 100,000 1,000,315 1,818
ER_100k_1m_E 100,000 1,000,044 1,666
ER_100k_1m_F 100,000 1,000,406 1,785
ER_100k_1m_G 100,000 1,001,171 1,724
ER_100k_1m_H 100,000 1,001,551 1,886
ER_100k_1m_I 100,000 1,000,237 1,639
ER_100k_1m_J 100,000 1,001,533 1,851
ER_100k_20m_F 100,000 20,001,732 107
ER_100k_20m_I 100,000 20,006,442 109
ER_100k_20m_J 100,000 20,003,479 109
ER_100k_4m_A 100,000 4,998,205 395
ER_100k_4m_C 100,000 4,999,271 398
ER_100k_4m_G 100,000 4,999,358 401
ER_100k_4m_J 100,000 4,996,501 414
ER_100k_5m_B 100,000 5,006,107 411
ER_100k_5m_D 100,000 5,001,575 404
ER_100k_5m_E 100,000 5,004,251 400
ER_100k_5m_F 100,000 5,002,190 400
ER_100k_5m_H 100,000 5,000,573 409
ER_100k_5m_I 100,000 5,001,846 400
ER_100k_999k_D 100,000 999,915 1,818
{longtblr}
[ theme = ams-theme, caption = Matrices and statistics in the narrow bandwidth data set used for the evaluation. The average wavefront size (Avg. wf) has been rounded down., label = table:random-bandwidth-graph-table, ] colspec = l | r | r | r , row1 = font=, rowhead = 1, rowfoot = 0 Matrix Size #Non-zeroes Avg. wf
NB_p14_b10_100k_A 100,000 146,565 87
NB_p14_b10_100k_B 100,000 146,328 115
NB_p14_b10_100k_C 100,000 147,201 61
NB_p14_b10_100k_D 100,000 146,972 73
NB_p14_b10_100k_E 100,000 147,369 73
NB_p14_b10_100k_F 100,000 146,855 111
NB_p14_b10_100k_G 100,000 147,350 132
NB_p14_b10_100k_H 100,000 147,412 85
NB_p14_b10_100k_I 100,000 147,132 132
NB_p14_b10_100k_J 100,000 146,781 105
NB_p3_b42_100k_A 100,000 127,045 46
NB_p3_b42_100k_B 100,000 127,019 55
NB_p3_b42_100k_C 100,000 127,708 29
NB_p3_b42_100k_D 100,000 127,341 45
NB_p3_b42_100k_E 100,000 127,569 67
NB_p3_b42_100k_F 100,000 127,137 47
NB_p3_b42_100k_G 100,000 127,774 52
NB_p3_b42_100k_H 100,000 127,029 46
NB_p3_b42_100k_I 100,000 127,475 39
NB_p3_b42_100k_J 100,000 127,275 62
NB_p5_b20_100k_A 100,000 102,053 1,298
NB_p5_b20_100k_B 100,000 102,621 1,063
NB_p5_b20_100k_C 100,000 102,021 1,298
NB_p5_b20_100k_D 100,000 102,968 1,075
NB_p5_b20_100k_E 100,000 102,650 952
NB_p5_b20_100k_F 100,000 102,309 1,162
NB_p5_b20_100k_G 100,000 103,152 892
NB_p5_b20_100k_H 100,000 102,324 1,190
NB_p5_b20_100k_I 100,000 102,465 1,369
NB_p5_b20_100k_J 100,000 102,244 1,010
Appendix B Time and space complexity of GrowLocal
Below we provide a more detailed discussion of the time and space complexity of GrowLocal. We first restate Theorem 3.1 more formally, including the assumption that the out-degrees and compute weights are within a constant factor.
[…]
3.1
Theorem 1 (formal).
Assume that there exist positive constants $\eta$ and $\varrho$ such that for all vertices $u,v\in V$, we have
$\omega(u)$ $\leq\eta\cdot\omega(v),$ (B.1)
$\rm{deg}^{+}(u)$ $\leq\varrho\cdot|E|/|V|.$ (B.2)
Then, the time complexity of GrowLocal is $O(|E|\cdot\log|V|)$ and the space complexity is $O(|E|)$.
Proof.
For each iteration $i$, denote by $\mathcal{V}_{p}^{(i)}$ the vertices which have been assigned to core $p$ in iteration $i$, by $\Gamma_{p}^{(i)}=|\mathcal{V}_{p}^{(i)}|$ the number of vertices assigned to core $p$ in iteration $i$, and by $\Omega_{p}^{(i)}$ the sum of weights of these vertices. From the algorithm design, specifically Line 3.1, we have for each iteration $i$ and each core $p$ that
$\Omega_{p}^{(i)}\leq\mu\Omega_{1}^{(i)},$ (B.3)
for some fixed positive constant $\mu$. We furthermore use the notations
$\Gamma_{\max}^{(i)}$ $=\max_{p}\Gamma_{p}^{(i)},$ (B.4)
$\Gamma_{\Sigma}^{(i)}$ $=\sum_{p}\Gamma_{p}^{(i)},\quad\text{and}$ (B.5)
$\omega_{\min}$ $=\min_{v\in V}\omega(v)$ (B.6)
for simplicity.
The key observation of the analysis is that during the formation of a superstep, the total number of vertices assigned over all the iterations is only a linear factor away from the number of vertices assigned in the final superstep. If we restrict ourselves to the first core, this is easy to see intuitively: the first iteration assigns $20$ vertices to core, the next iteration assigns $20\cdot\tfrac{3}{2}$, the following assigns $20\cdot(\tfrac{3}{2})^{2}$, and so on. If the iteration that is accepted in the end assigns $\alpha^{*}$ vertices to the first core, then the preceding iterations assign at most $\alpha^{*}\sum_{i=1}^{\infty}(\frac{2}{3})^{i}$ altogether, and the last examined iteration (which is rejected) possibly also assigns $\frac{3}{2}\alpha^{*}$; this is altogether still in $O(\alpha^{*})$. Note that the ratio between the last two iterations can also be less than $\frac{3}{2}$, but this does not affect the claim.
Extending the argument above to all the cores is slightly more technical due to the different vertex weights. For a core $p$, we have
$\Gamma_{p}^{(i)}=\sum_{v\in\mathcal{V}_{p}^{(i)}}1\leq\sum_{v\in\mathcal{V}_{p}^{(i)}}\eta\frac{\omega(v)}{\sum_{u\in\mathcal{V}_{1}^{(i)}}\omega(u)/\sum_{u\in\mathcal{V}_{1}^{(i)}}1}\\ =\eta\frac{\Omega_{p}^{(i)}}{\Omega_{1}^{(i)}}\sum_{u\in\mathcal{V}_{1}^{(i)}}1\leq\eta\mu\sum_{u\in\mathcal{V}_{1}^{(i)}}1=\eta\mu\Gamma_{1}^{(i)}.$ (B.7)
Here, we used (B.1) and (B.3). Thus, if $\alpha^{(i)}=20\cdot(\tfrac{3}{2})^{i-1}$ is the parameter used for iteration $i$, then we have
$\alpha^{(i)}\leq\Gamma_{\max}^{(i)}\leq\eta\mu\cdot\alpha^{(i)}.$ (B.8)
This means that for iterations $i$ and $j$ with $i<j$, we get that the ratio $\Gamma_{\max}^{(i)}/\Gamma_{\max}^{(j)}$ is at most $\eta\mu\cdot(\frac{2}{3})^{j-i}$.
For the proof below, we actually require a similar upper bound on the ratio
$\frac{\Gamma_{\max}^{(i)}+L}{\Gamma_{\max}^{(j)}+L}$ (B.9)
instead. For this, we separate two cases, namely the first few iterations and all the remaining ones. Recall that $L$ was chosen as a constant. Assume first that we have $\Gamma_{\max}^{(i)}\geqslant L$. In this case, we can upper bound the expression above by
$\frac{2\cdot\Gamma_{\max}^{(i)}}{\Gamma_{\max}^{(j)}}\leqslant 2\eta\mu\cdot\left(\tfrac{2}{3}\right)^{j-i}\,.$ (B.10)
On the other hand, assume that $\Gamma_{\max}^{(i)}<L$. In this case, we can upper bound (B.9) simply by $1$. The key observation is that the number of these iterations with $\Gamma_{\max}^{(i)}<L$ is at most a constant in any superstep. Indeed, starting with $\alpha=20$ and multiplying by $\frac{3}{2}$ each round, we already have $\alpha>L$ by the $\lceil\log(L)/\log(1.5)\rceil$-th iteration, and hence $\Gamma_{\max}^{(i)}\geqslant L$ for $i\geq C_{L}\coloneqq\lceil\log(L)/\log(1.5)\rceil$. As such, there are only $C_{L}=O(\log(L))=O(1)$ distinct values where we will use this upper bound of $1$.
Our algorithm only accepts and saves an iteration if its parallelization rate $\beta$ is at least a $0.97$ factor of the best parallelization rate observed so far during this superstep. Hence, if iteration $j$ is accepted and $i$ is any iteration such that $i<j$, then we have
$\frac{\sum_{p}\Omega_{p}^{(j)}}{\max_{p}\Omega_{p}^{(j)}+L}\geqslant 0.97\cdot\frac{\sum_{p}\Omega_{p}^{(i)}}{\max_{p}\Omega_{p}^{(i)}+L}\,.$ (B.11)
Due to our assumption on the vertex weights, we have for any iteration $i$ that
$\omega_{\min}\cdot\Gamma_{\max}^{(i)}\leqslant\max_{p}\Omega_{p}^{(j)}\leqslant\eta\omega_{\min}\cdot\Gamma_{\max}^{(i)}$ (B.12)
as well as
$\omega_{\min}\cdot\Gamma_{\Sigma}^{(i)}\leqslant\sum_{p}\Omega_{p}^{(i)}\leqslant\eta\omega_{\min}\cdot\Gamma_{\Sigma}^{(i)}.$ (B.13)
Using these in Inequality (B.11), we get that
$\frac{\eta\omega_{\min}\cdot\Gamma_{\Sigma}^{(j)}}{\omega_{\min}\cdot\Gamma_{\max}^{(j)}+L}\geqslant 0.97\cdot\frac{\omega_{\min}\cdot\Gamma_{\Sigma}^{(i)}}{\eta\omega_{\min}\cdot\Gamma_{\max}^{(i)}+L}\,,$ (B.14)
which further implies
$\Gamma_{\Sigma}^{(i)}\leqslant\frac{\eta^{2}}{0.97}\cdot\frac{\Gamma_{\max}^{(i)}+L}{\Gamma_{\max}^{(j)}+L}\cdot\Gamma_{\Sigma}^{(j)}\,.$ (B.15)
Using our upper bounds on (B.9) and $\eta=O(1)$, this implies
$\Gamma_{\Sigma}^{(i)}\leqslant O(1)\cdot\Gamma_{\Sigma}^{(j)}\,$ (B.16)
for $i\in\{1,...,C_{L}-1\}$, and
$\Gamma_{\Sigma}^{(i)}\leqslant O(1)\cdot\left(\tfrac{2}{3}\right)^{j-i}\cdot\Gamma_{\Sigma}^{(j)}\,$ (B.17)
for $i\geqslant C_{L}$. Assuming that iteration $j$ is the final worthy iteration that is accepted for our superstep, this means that the total number of assigned vertices made in iterations $i\in\{1,2,...,j-1\}$ is at most
$\sum_{i=1}^{j-1}\Gamma_{\Sigma}^{(i)}\leqslant O(1)\cdot\Gamma_{\Sigma}^{(j)}\cdot\left(O(1)+\sum_{\ell=1}^{j-1}\left(\tfrac{2}{3}\right)^{\ell}\right)\,.$ (B.18)
With the geometric sum upper bounded by $2$, we get that the number of assigned vertices is indeed in $O(\Gamma_{\Sigma}^{(j)})$.
Note that the argument above does not consider the possible last iteration $(j+1)$ which is rejected by our algorithm. Nevertheless, we can bound the assignments here with a similar argument. If the iteration was rejected, then its parallelization score is at most as high as that of iteration $j$, i.e., $\beta^{(j+1)}\leqslant\beta^{(j)}$. As before, this implies
$\frac{\omega_{\min}\cdot\Gamma_{\Sigma}^{(j+1)}}{\omega_{\min}\cdot\eta\cdot\Gamma_{\max}^{(j+1)}+L}\leqslant\frac{\omega_{\min}\cdot\eta\cdot\Gamma_{\Sigma}^{(j)}}{\omega_{\min}\cdot\Gamma_{\max}^{(j)}+L}\,,$ (B.19)
and hence
$\Gamma_{\Sigma}^{(j+1)}\leqslant O(1)\cdot\frac{\Gamma_{\max}^{(j+1)}+L}{\Gamma_{\max}^{(j)}+L}\cdot\Gamma_{\Sigma}^{(j)}\,.$ (B.20)
As
$\frac{\Gamma_{\max}^{(j+1)}+L}{\Gamma_{\max}^{(j)}+L}$ $=\frac{\Gamma_{\max}^{(j+1)}-\Gamma_{\max}^{(j)}}{\Gamma_{\max}^{(j)}+L}+1$ (B.21)
$\leqslant\frac{\Gamma_{\max}^{(j+1)}}{\Gamma_{\max}^{(j)}}+1\leqslant\tfrac{3}{2}\eta\mu+1\,,$
we have that $\Gamma_{\Sigma}^{(j+1)}$ is again in $O(\Gamma_{\Sigma}^{(j)})$.
As such, the number of assignments in any superstep is linear in the size of the vertices that are finally scheduled. Summing this up over all the supersteps, we get that the algorithm altogether only makes $O(|V|)$ assignments over all supersteps and iterations.
For each assignment, the chosen vertex is selected from a priority queue data structure. Each such data structure contains at most $|V|$ vertices, so the cost of each assignment is $O(\log|V|)$. However, after each assignment of a concrete vertex $v$, we also need to examine all the children $u$ of $v$, check if $u$ also becomes ready with this assignment (i.e., all parents of $u$ are computed now), and if so, then also insert $u$ into such a priority queue at a time cost of $O(\log|V|)$. Since any vertex $v$ has at most $\varrho\cdot|E|/|V|$ children and $\varrho\in O(1)$, this sums up to a total of $O(|V|\cdot|E|/|V|\cdot\log|V|)$ over all the $O(|V|)$ assignments. This results in an overall time complexity of $O(|E|\cdot\log|V|)$ for the algorithm.
The space complexity of the algorithm is much easier to settle: each iteration only stores $O(|V|)$ data, and we store at most two iterations at a time, so the main bottleneck here is simply storing the input DAG itself, which requires $O(|E|)$ space. ∎
For the sake of completeness, we complement the theoretical bound in Theorem 3.1 with empirical data in Figure B.1.
Figure B.1. Scheduling time of Funnel+GL and GrowLocal on the SuiteSparse data set. The straight lines are the best square-mean-error fit of the family of curves $\log(y)=\log(x)+c$, where $c$ is a free parameter.
Appendix C Further discussion
C.1. Comparison to barrier list schedulers
Recall that one of the recent works on DAG scheduling with synchronization barriers is that of Papp et al.  [PAKY24], which analyzes schedulers not for a concrete application, more abstractly in terms of BSP cost. While the so-called BSPg scheduling heuristic in this work is rather different from our algorithm, the idea of prioritizing vertices in GrowLocal that are computable exclusively on a specific core was inspired by this algorithm. To show for completeness that our scheduler also significantly outperforms this BSPg algorithm, we also ran this BSPg scheduling algorithm as a baseline. The results show that GrowLocal achieves a factor 8.31×8.31$\times$8.31 × geometric-mean speed-up to BSPg on the SutieSparse data set.
C.2. On the synchronisation parameter $L$
Recall that the parameter $L$ in GrowLocal, Algorithm 3.1, represents the time cost of inserting a synchronization barrier, and is used to determine the parallelization rate in our GrowLocal algorithm.
If we consider the compute time of basic operations (additions or multiplications) with double precision numbers, and compare this to the time of synchronization, we get that the correct magnitude of $L$ ranges from a few hundreds to a few thousands on modern computing architectures. We ran some preliminary experiments with a few different choices of $L$ on this order of magnitude, and chose a value of $L=500$ based on these empirical observations.